Circle Circumfe Circle Diameter r is radius Finish A sprinter runs the curve of this 200 m in 10.34 s. Assume he ran in a lane which makes a semicircle (r 36.3 m) for the first part of the race. At 3.44 s into the race his speed is 4.0 m/s. At 8.0 s into the race his speed is 9.6 m/s. His speed after the curve was 12.8 m/s. What was the sprinters radial acceleration at 3.44 s? Answer: Check
Answer #1

As the spinter is along the curvature so the radial acceleration acts , towards the center which is given by

a_rad = v^2/r

v is speed, r is radius of the curve

given v= 4 m/s , r = 36.3 m

a_rad = 4^2/36.3 m/s^2

a_rad= 0.441 m/s^2